OpenStudy (anonymous):

Let A(-7,4) and B(5,-12) be points on a plane. Find the: a) length of segment b) equation of the perpendicular bisector of AB c) equation of the circle for which AB is a diameter

OpenStudy (anonymous):

Length of segment= sqrt{(5+7)^2 + (-12-4)^2} =sqrt{144+256) =20 unit

OpenStudy (anonymous):

(a) Length is given by the Pythagorean theorem:\[\sqrt{(5-(-7))^2+(-12-4)^2}\](b) Slope is the inverse reciprocal of the slope of the line connecting the two points and will intersect the midpoint between the two points. Thus, by point-slope form, which is left for you to simplify:\[y-\left(\dfrac{4+(-12)}{2}\right)=-\dfrac{5-(-7)}{-12-4}\left[x-\left(\dfrac{-7+5}{2}\right)\right]\](c) Given by the transformation of the generic formula for a circle given that the midpoint of the two given points is the center of the circle and the radius is half the length of the line connecting them, left for you to simplify:\[\left[\dfrac{\sqrt{(5-(-7))^2+(-12-4)^2}}{2}\right]^2=\left[y-\left(\dfrac{4+(-12)}{2}\right)\right]^2+\left[x-\left(\dfrac{-7+5}{2}\right)\right]^2\]

OpenStudy (anonymous):

eqn of perpendicular bisector --- first slope of line AB =- (-12-4)/(5+7) =4/3 then slope perpendicular bisector of AB=-3/4 and mid pt of line AB={(-7+5)/2 , (4-12)/2} =(-1,-4) so eqn of perpendicular bisector of AB will be y+4=-3/4(x+1) {using y-y1=m(x-x1)} 4y+4=-3x-3 => 3x+4y + 7=0

OpenStudy (anonymous):

it says teh asnwer is 3x-4y = 13

OpenStudy (anonymous):

for teh equation of the perpemdicular bisector i mean

OpenStudy (anonymous):

did you get the length of the line segment?

OpenStudy (anonymous):

yeah i understand taht...its 20 with the distance formula. i'm stuck on the otehr 2 now

OpenStudy (anonymous):

\[\sqrt{(5-(-7))^2+(-12-4)^2}=\sqrt{12^2+16^2}=20\]

OpenStudy (anonymous):

thanks taht shows all teh steps

OpenStudy (anonymous):

you need the midpoint right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[(\frac{-7+5}{2},\frac{4-12}{2})\] \[(-1,-4)\]

OpenStudy (anonymous):

is that ok? average of the coordinates

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

ok now the equation for the circle. center is \[(-1,-4)\] radius is 10 (half of 20) so equation for the circle is \[(x+1)^2+(y+4)^2=10^2\]

OpenStudy (anonymous):

or if you prefer \[(x+1)^2+(y+4)^2=100\]

OpenStudy (anonymous):

now for the equation of the perpendicular bisector

OpenStudy (anonymous):

(à la what I wrote above)

OpenStudy (anonymous):

first we need the slope of the line between the two points \[(-7,4) , (5,-12)\] it is \[m=\frac{-12-4}{5+7}=\frac{-16}{12}=-\frac{4}{3}\]

OpenStudy (anonymous):

so the slope of the perpendicular line is \[\frac{3}{4}\] clear so far?

OpenStudy (anonymous):

yeah its negative recipricol

OpenStudy (anonymous):

ok and so we have the slope, and also the point, namely \[(-1,-4)\] so use the all mighty point slope formula \[y-y_1=m(x-x_1)\]so get the equation. it is \[y+4=\frac{3}{4}(x+1)\]

OpenStudy (anonymous):

now if you want to write this in some other form, say slope - intercept for or standard form you can change it however you like. for example you could write \[y+4=\frac{3}{4}x+\frac{3}{4}\] \[y=\frac{3}{4}x-\frac{13}{4}\]

OpenStudy (anonymous):

how did they get teh answer 3x-4y = 13? what format is that

OpenStudy (anonymous):

dumb math teacher form, but we can get there too

OpenStudy (anonymous):

lol..yupp math teachers liek to make everything look and sound complicated

OpenStudy (anonymous):

\[y=\frac{3}{4}x-\frac{13}{4}\] \[4y=3x-13\] \[13=3x-4y\] whatever you like

OpenStudy (anonymous):

\[3x-4y=13\] \[3x-4y-13=0\] anything

OpenStudy (anonymous):

wasn't that bad was it?

OpenStudy (anonymous):

no it actually wasnt. =) except when see teh question at first it drives me crazy

OpenStudy (anonymous):

i have a question. For the equation 9x^2 + 16y^2 = 144 is it posible to make it into y=mx+b form

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

so how wud u draw a graph for that

OpenStudy (anonymous):

for one thing \[y=mx+b\] is a line and you have an ellipse! http://www.wolframalpha.com/input/?i=9x^2+%2B+16y^2+%3D+144

OpenStudy (anonymous):

oh i see...so how did tehy get those points or liek how did they draw the ellipse

OpenStudy (anonymous):

like for a circle u have a center n a radius so u can draw it...how about this? what do each of teh numbers mean

OpenStudy (anonymous):

there is no snap answer to this question. you can write it in the form \[\frac{x^2}{16}+\frac{y^2}{9}=1\] and then look at the standard form on the ellipse. http://mysite.du.edu/~jcalvert/math/ellipse.htm

OpenStudy (anonymous):

do u sqt root the values of 9 and 16 b.c in teh drawing on teh x values it reaches close to 4 and on teh y value it is close to 3

OpenStudy (anonymous):

im confused how u sketch a ellipse with teh equation 9x^2 + 16y^2 = 144

OpenStudy (anonymous):

@august151985....r u there